Find the interest rate (or rates of return) in each of the following situations. Do not round intermediate calculations. Round your answers to the nearest whole number.

  1. You borrow $650 and promise to pay back $676 at the end of 1 year.
  2.   %
  3. You lend $650 and receive a promise to be paid $676 at the end of 1 year.
  4.   %
  5. You borrow $60,000 and promise to pay back $289,075 at the end of 12 years.
  6.   %
  7. You borrow $16,000 and promise to make payments of $5,001.00 at the end of each of the next 5 years.
1
Expert's answer
2021-03-10T02:54:07-0500

Let Future Value = FV

Present Value = PV

Number of Years = n

Interest Rate = r

"FV = PV(1+r)" n

"r = (FV\/PV)"1/n -1

(a)

"PV = 650"

"FV = 676"

"n = 1"

"r = (676\/650)" 1/1"- 1 = 0.04 or 4%" %

(b)

"PV = 650"

"FV = 676"

"n = 1"

"r = (676\/650)" 1/1 "- 1 = 0.04 or 4%" %

(c)

"PV = 60000"

"FV = 289075"

"n = 12"

"r = (289075\/60000)" 1/12 "- 1 = 0.14 or 14%" %

(d)

"PV = 16000"

Payment received at end of each year=5001

"n = 5"

"=16000 = 5001[1- (1+r)1\/5]\/r" 

"=17" %